# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v1

@@ A.4 Further results
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 9.7 0.6 0.3 9.2 0.4 0.2 10.5 0.4 0.2 12.0 0.5 0.1 13.3 0.5 0.0 13.7 0.5 0.3 11.6 0.9 0.3 12.3 0.8 0.2 12.3 0.9 0.3 11.7 0.7 0.3 11.2 0.8 0.6 13.9 0.4 1.3
BeerRNN 8.2 0.4 3.6 12.1 0.5 0.8 9.3 0.4 1.3 13.3 0.4 1.6 10.9 0.3 1.3 13.1 0.5 1.5 12.4 0.9 1.7 12.7 0.8 2.1 11.9 0.9 2.0 11.7 0.8 1.9 11.4 0.8 1.1 14.2 0.4 2.5
Bouali 4.7 0.1 0.7 7.0 0.3 0.5 12.1 0.2 0.6 11.6 0.1 0.6 13.1 0.2 0.3 13.5 0.4 1.5 11.2 0.7 2.2 12.0 0.6 3.4 12.3 0.5 2.2 9.8 0.5 1.5 12.9 0.7 2.6 14.0 0.3 2.5
Burke Shaw 3.4 0.1 0.2 8.0 0.1 0.3 5.6 0.1 0.2 5.1 0.2 0.1 11.1 0.2 0.4 13.2 0.6 1.0 11.0 0.6 0.9 11.0 0.7 0.5 11.7 0.8 0.2 10.0 0.7 1.1 10.2 0.7 0.6 12.9 0.4 0.8
Ca Two Plus Quasip. 3.3 0.1 0.3 9.6 0.5 1.2 9.5 0.3 0.3 12.6 0.2 0.9 12.7 0.4 0.6 13.2 0.4 1.6 9.4 0.7 0.8 9.4 0.6 1.5 9.6 0.7 0.8 11.5 0.5 0.9 11.1 0.6 2.0 13.8 0.2 5.2
Chen Lee 5.1 0.4 0.1 4.2 0.2 0.0 3.6 0.2 0.0 3.9 0.2 0.0 5.3 0.3 0.0 12.0 0.8 0.3 4.5 0.8 0.3 8.8 0.7 1.1 7.7 0.7 0.5 7.2 0.7 0.5 11.3 0.8 0.2 8.9 0.4 0.6
Chua 3.0 0.1 0.4 12.9 0.2 0.4 12.4 0.4 0.4 13.3 0.2 0.1 13.5 0.4 0.4 13.5 0.3 0.9 11.9 0.6 1.0 11.7 0.4 0.6 11.7 0.4 0.6 12.0 0.4 0.7 12.1 0.4 0.2 13.4 0.3 1.0
Coullet 0.8 0.0 0.3 0.6 0.0 0.1 6.3 0.0 0.1 3.9 0.1 0.1 13.7 0.3 0.1 14.6 0.3 0.6 11.4 0.7 0.6 12.8 0.7 0.4 11.8 0.6 0.7 10.7 0.6 1.1 13.3 0.6 0.7 15.0 0.2 1.8
Dadras 3.0 0.2 13.9 10.0 0.5 9.2 9.2 0.5 4.2 11.8 0.4 7.5 11.4 0.3 4.7 12.9 0.8 9.0 10.5 0.8 8.9 9.4 0.7 9.0 11.0 0.6 8.8 8.8 0.5 8.9 13.1 0.8 7.9 12.2 0.4 9.2
Dequan Li 10.0 0.5 0.1 12.6 0.5 0.1 13.2 0.4 0.1 13.1 0.3 0.1 12.9 0.4 0.5 13.1 0.3 0.4 12.4 0.4 0.4 12.7 0.4 0.3 12.2 0.4 0.3 12.6 0.4 0.6 12.3 0.4 1.1 13.0 0.4 0.4
Finance 3.0 0.2 0.6 8.5 0.4 0.2 9.4 0.6 0.3 13.4 0.3 0.5 12.2 0.4 0.3 13.6 0.5 0.3 12.0 0.7 1.3 11.7 0.6 2.3 12.4 0.7 1.3 11.8 0.6 1.4 12.6 0.8 2.3 14.0 0.3 2.1
Genesio Tesi 10.5 0.2 0.3 8.9 0.1 0.1 8.7 0.1 0.0 7.7 0.1 0.0 11.8 0.4 0.2 12.3 0.3 0.3 10.7 0.3 0.2 11.4 0.9 0.8 10.6 0.6 1.1 10.0 0.4 1.6 10.4 0.8 0.4 12.3 0.2 0.7
Guckenheimer Holmes 7.9 0.0 0.2 10.6 0.4 0.1 11.7 0.2 0.0 11.3 0.4 0.1 13.2 0.7 0.1 12.5 0.2 0.1 10.7 0.6 0.1 10.6 0.6 0.2 9.2 0.6 0.1 10.3 0.6 0.1 10.4 0.5 0.5 11.6 0.3 0.6
Hadley 8.4 0.3 0.2 10.8 0.5 0.2 9.3 0.6 0.2 11.9 0.4 0.2 12.5 0.4 0.2 12.5 0.3 0.4 10.9 0.7 0.4 10.7 0.7 1.0 9.5 0.6 0.3 10.2 0.4 0.4 11.7 0.7 0.8 12.5 0.3 1.6
Halvorsen 2.8 0.1 0.2 6.0 0.1 0.2 8.4 0.1 0.2 5.9 0.1 0.3 12.6 0.2 0.3 13.3 0.3 1.3 11.8 0.8 1.1 9.5 0.7 0.8 8.9 0.8 0.4 10.4 0.5 0.4 12.2 0.6 1.3 13.2 0.3 1.5
Hindmarsh Rose 6.8 0.5 0.6 9.7 0.3 0.1 6.7 0.5 0.3 8.4 0.1 0.1 11.5 0.4 0.7 13.8 0.6 1.0 12.7 0.7 1.3 12.5 0.7 1.5 13.9 0.6 1.1 13.4 0.7 1.1 13.9 0.6 0.7 14.0 0.2 1.6
Isothermal Chemical 2.5 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.0 13.8 0.5 0.2 7.9 0.8 0.3 5.9 0.8 0.2 8.4 0.8 0.2 8.5 0.8 0.2 12.8 0.8 0.5 15.6 0.2 2.8
Itik Banks tumor 9.1 0.2 1.2 10.0 0.6 0.5 8.4 0.5 0.3 10.5 0.4 0.3 12.8 0.3 0.2 12.1 0.7 0.4 9.2 0.7 1.2 9.5 0.7 1.2 9.7 0.8 1.2 9.6 0.7 1.0 13.3 0.9 0.5 13.2 0.3 1.7
Jerk Circuit 2.7 0.3 0.4 10.1 0.4 1.0 9.7 0.4 0.4 7.5 0.3 0.4 11.1 0.4 0.4 12.3 0.3 0.2 10.8 0.8 1.6 10.9 0.5 2.0 10.5 0.9 1.9 10.1 0.9 1.8 11.1 0.9 1.8 13.9 0.4 2.1
Laser 6.5 0.3 0.1 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.2 13.8 0.5 0.4 14.1 0.5 0.3 11.8 0.7 0.3 11.5 0.7 0.4 11.0 0.7 0.5 11.9 0.7 1.6 13.3 0.6 0.7 13.8 0.2 1.2
Lorenz 1.1 0.1 0.7 6.6 0.2 0.3 6.0 0.2 0.3 10.0 0.2 0.2 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.3 11.3 0.5 0.6 12.7 0.6 0.9 12.8 0.3 1.8 11.4 0.7 0.2 13.0 0.3 1.9
Lorenz 84 6.9 0.3 4.9 9.5 0.4 5.2 11.0 0.4 4.8 11.5 0.4 3.2 12.2 0.4 2.5 12.9 0.5 4.4 12.9 0.8 2.4 12.9 0.8 2.5 11.5 0.9 2.4 12.3 0.7 2.4 13.4 0.8 3.3 13.4 0.5 2.4
Lu Chen 3.4 0.5 0.8 6.3 0.4 0.2 5.3 0.3 0.3 8.0 0.1 0.2 7.1 0.1 0.2 4.9 0.2 0.1 11.6 0.9 1.3 11.9 1.0 1.2 11.3 1.0 1.9 11.4 1.0 1.4 9.7 0.8 0.7 13.7 0.4 3.4
Moore Spiegel 5.9 0.4 1.0 10.4 0.6 1.0 10.0 0.5 0.0 12.5 0.4 0.2 9.9 0.5 0.6 12.1 0.3 0.9 10.5 0.9 2.8 11.4 0.9 1.8 10.1 0.9 2.5 9.8 0.7 3.4 9.4 0.8 4.6 12.5 0.3 3.6
Multi Chua 7.8 0.3 0.8 9.3 0.5 0.0 9.3 0.2 0.2 7.1 0.4 0.1 13.3 0.2 0.4 14.0 0.3 1.2 11.5 0.6 1.8 12.2 0.4 3.1 9.5 0.5 2.6 10.1 0.4 2.4 10.6 0.4 1.2 14.0 0.3 2.8
Nose Hoover 10.0 0.3 0.3 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.5 13.0 0.5 0.5 10.6 0.9 0.8 10.6 0.9 0.7 10.1 0.7 0.7 10.2 0.5 0.4 10.2 0.7 0.5 12.9 0.4 0.4
Pehlivan Wei 4.0 0.2 0.3 8.5 0.4 0.4 10.1 0.4 0.3 11.3 0.3 0.3 10.8 0.4 0.3 12.6 0.5 0.7 9.2 0.7 1.0 10.1 0.8 1.0 9.2 0.7 0.6 10.6 0.8 1.3 9.3 0.7 1.0 12.2 0.6 1.9
Sakarya 1.5 0.2 0.3 5.6 0.5 0.2 5.0 0.3 0.2 7.5 0.2 0.3 11.1 0.4 0.2 12.5 0.4 0.4 10.2 0.9 0.4 10.0 0.6 1.4 7.7 0.9 0.9 9.1 0.4 0.8 13.1 0.8 0.8 12.6 0.4 0.7
San Um Srisuchinwong 4.1 0.2 0.2 12.9 0.4 1.0 12.7 0.5 1.6 11.4 0.5 1.0 13.3 0.5 1.2 13.0 0.6 1.1 12.7 0.7 0.9 10.4 0.7 0.9 13.4 0.6 1.2 11.7 0.7 1.5 10.7 0.7 1.6 13.1 0.4 2.4
Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.2 11.8 0.3 0.3 12.3 0.3 1.1 10.5 0.8 1.0 8.1 0.5 1.6 10.0 0.7 0.7 10.4 0.7 1.0 9.9 0.7 0.8 12.0 0.3 3.0
Sprott B 2.4 0.2 0.3 4.3 0.2 0.8 5.6 0.4 1.6 11.7 0.5 0.2 12.7 0.4 0.6 13.2 0.7 1.7 11.8 0.8 0.8 11.5 0.6 0.6 12.3 0.8 1.2 10.6 0.7 1.4 12.5 0.8 1.8 12.6 0.4 1.5
Sprott C 2.3 0.1 0.6 11.6 0.4 1.2 12.4 0.4 0.7 11.9 0.4 1.2 13.3 0.3 0.3 13.1 0.4 2.4 10.0 0.5 4.7 9.2 0.5 3.9 10.7 0.5 5.3 9.3 0.5 4.3 12.8 0.6 2.9 13.3 0.2 2.4
Sprott D 4.0 0.2 0.3 9.8 0.6 0.1 10.3 0.4 0.1 12.9 0.3 0.2 13.4 0.5 0.4 13.4 0.4 0.1 12.2 0.7 0.2 10.9 0.6 0.2 9.9 0.5 0.6 11.5 0.6 0.6 12.9 0.7 0.2 14.0 0.2 1.4
Sprott F 3.0 0.2 0.3 9.6 0.5 0.3 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 0.7 12.7 0.4 0.6 9.9 0.7 1.5 10.4 0.7 1.3 10.4 0.6 1.0 12.1 0.5 1.1 11.4 0.8 1.1 13.3 0.3 0.9
Sprott G 2.3 0.1 0.7 7.0 0.6 0.1 7.7 0.5 0.1 10.0 0.3 0.3 7.2 0.1 0.4 13.0 0.3 0.6 10.9 0.9 1.2 11.4 0.6 1.3 9.2 0.8 1.4 10.8 0.8 1.7 14.1 0.8 0.7 13.1 0.2 1.3
Sprott I 2.7 0.1 0.5 9.7 0.5 0.4 11.3 0.2 0.4 9.0 0.1 0.4 11.9 0.3 0.5 13.7 0.2 1.1 10.2 0.9 2.1 9.8 0.9 3.4 8.4 0.6 2.9 9.1 0.6 2.8 12.2 0.7 1.9 14.4 0.2 6.2
Sprott L 3.4 0.1 0.4 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.2 13.0 0.2 0.1 10.5 0.9 1.2 10.4 0.8 1.8 11.4 0.9 1.2 12.1 0.7 1.0 10.8 0.8 1.3 14.2 0.3 3.1
Sprott M 3.8 0.1 0.4 9.8 0.4 0.0 10.5 0.4 0.1 12.4 0.1 0.1 12.9 0.2 0.2 14.0 0.6 0.6 9.8 0.8 1.4 10.3 0.6 1.6 9.9 0.8 1.5 11.8 0.6 1.6 11.6 0.7 1.5 14.3 0.2 3.2
Sprott O 1.3 0.1 0.9 3.6 0.1 0.3 4.6 0.0 0.5 9.7 0.1 0.4 12.4 0.4 0.4 12.9 0.4 0.9 10.8 1.0 1.3 12.6 0.9 1.3 10.4 0.8 1.6 11.5 0.9 1.1 10.7 0.8 0.6 14.0 0.4 2.0
Sprott P 5.5 0.2 0.6 10.8 0.3 2.2 11.9 0.3 1.3 12.7 0.3 1.1 12.7 0.3 1.2 13.3 0.3 2.0 7.3 0.5 3.1 10.6 0.6 4.5 8.3 0.6 4.2 10.0 0.6 5.0 11.1 0.7 2.7 12.2 0.3 5.7
Sprott Q 4.5 0.1 0.5 10.1 0.2 0.2 7.6 0.1 0.2 7.3 0.1 0.3 11.6 0.2 0.2 13.2 0.3 0.8 11.8 0.8 2.5 11.6 0.7 1.1 11.9 0.6 0.5 11.2 0.6 2.8 10.9 0.7 0.7 13.1 0.4 4.3
Sprott R 3.5 0.3 0.4 9.3 0.8 0.1 6.7 0.5 0.2 11.0 0.5 0.4 12.6 0.5 0.6 13.1 0.4 0.3 10.6 0.8 0.4 10.4 0.5 0.3 9.6 0.7 0.8 10.5 0.7 0.8 11.8 0.8 0.8 13.8 0.3 1.3
Sprott S 5.3 0.1 0.6 10.1 0.9 1.7 9.9 0.6 0.7 12.3 0.2 0.9 12.6 0.4 0.4 13.1 0.3 1.1 10.6 0.7 0.6 9.7 0.8 3.5 11.0 0.6 1.8 10.4 0.4 3.7 11.8 0.8 1.8 11.6 0.3 4.6
Sprott Torus 3.2 0.2 0.1 5.4 0.7 0.4 10.4 0.7 0.1 14.3 0.6 0.2 14.5 0.6 0.4 13.8 0.7 0.3 6.2 0.8 2.9 8.3 0.6 5.2 5.3 0.8 5.4 5.8 0.7 4.1 13.6 0.8 1.0 15.5 0.3 2.1
[…]

### v2

@@ A.7 Comparison of DynaMix to custom-trained DSR models
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 10.4 nan 0.0 9.2 0.4 0.0 10.5 0.4 0.0 12.0 0.5 0.0 13.3 0.5 0.0 13.7 0.5 0.1 11.6 0.9 0.1 12.3 0.8 0.2 12.3 0.9 0.1 11.7 0.7 0.2 11.2 0.8 0.7 13.9 0.4 1.6 11.8 0.4 0.1
BeerRNN 5.2 0.7 0.6 12.1 0.5 0.5 9.3 0.4 0.7 13.3 0.4 0.7 10.9 0.3 0.6 13.1 0.5 0.8 12.4 0.9 1.0 12.7 0.8 1.7 11.9 0.9 1.5 11.7 0.8 1.6 11.4 0.8 1.0 14.2 0.4 1.5 14.0 0.3 0.5
Bouali 5.9 0.1 0.2 7.0 0.3 0.4 12.1 0.2 0.5 11.6 0.1 0.5 13.1 0.2 0.4 13.5 0.4 2.0 11.2 0.7 2.3 12.0 0.6 3.1 12.3 0.5 2.0 9.8 0.5 2.2 12.9 0.7 4.4 14.0 0.3 5.6 9.7 0.4 1.1
Burke Shaw 3.8 0.2 0.3 8.0 0.1 0.5 5.6 0.1 0.4 5.1 0.2 0.2 11.1 0.2 0.5 13.2 0.6 1.1 11.0 0.6 1.3 11.0 0.7 0.9 11.7 0.8 0.9 10.0 0.7 1.2 10.2 0.7 1.2 12.9 0.4 2.1 12.4 0.3 0.8
Ca Two Plus Quasip. 2.8 0.1 0.3 9.6 0.5 1.6 9.5 0.3 0.5 12.6 0.2 1.2 12.7 0.4 1.4 13.2 0.4 1.1 9.4 0.7 2.6 9.4 0.6 1.3 9.6 0.7 1.4 11.5 0.5 2.4 11.1 0.6 4.0 13.8 0.2 4.6 13.4 0.3 1.9
Chen Lee 3.6 0.3 0.2 4.2 0.2 0.0 3.6 0.2 0.1 3.9 0.2 0.1 5.3 0.3 0.1 12.0 0.8 0.5 4.5 0.8 0.4 8.8 0.7 0.9 7.7 0.7 0.6 7.2 0.7 1.4 11.3 0.8 0.8 8.9 0.4 2.0 4.6 0.2 0.3
Chua 2.9 0.1 0.8 12.9 0.2 0.6 12.4 0.4 0.8 13.3 0.2 0.2 13.5 0.4 1.0 13.5 0.3 2.1 11.9 0.6 2.4 11.7 0.4 1.3 11.7 0.4 1.8 12.0 0.4 2.2 12.1 0.4 0.8 13.4 0.3 4.0 12.7 0.2 1.0
Coullet 0.9 0.0 0.2 0.6 0.0 0.3 6.3 0.0 0.2 3.9 0.1 0.2 13.7 0.3 0.2 14.6 0.3 0.8 11.4 0.7 1.1 12.8 0.7 1.5 11.8 0.6 1.1 10.7 0.6 3.0 13.3 0.6 1.5 15.0 0.2 3.5 9.3 0.2 0.5
Dadras 2.2 0.2 1.3 10.0 0.5 4.2 9.2 0.5 1.9 11.8 0.4 2.3 11.4 0.3 3.5 12.9 0.8 4.1 10.5 0.8 5.4 9.4 0.7 5.0 11.0 0.6 6.8 8.8 0.5 5.6 13.1 0.8 5.0 12.2 0.4 6.0 13.1 0.6 4.5
Dequan Li 8.5 0.1 0.8 12.6 0.5 0.2 13.2 0.4 0.3 13.1 0.3 0.3 12.9 0.4 1.4 13.1 0.3 1.0 12.4 0.4 0.8 12.7 0.4 0.8 12.2 0.4 0.6 12.6 0.4 1.4 12.3 0.4 3.3 13.0 0.4 1.7 13.0 0.3 1.3
Finance 3.1 0.1 0.3 8.5 0.4 0.7 9.4 0.6 0.5 13.4 0.3 0.6 12.2 0.4 0.9 13.6 0.5 1.1 12.0 0.7 2.8 11.7 0.6 4.0 12.4 0.7 3.7 11.8 0.6 4.7 12.6 0.8 4.9 14.0 0.3 4.2 13.2 0.5 2.7
Genesio Tesi 9.8 0.2 0.4 8.9 0.1 0.2 8.7 0.1 0.1 7.7 0.1 0.1 11.8 0.4 0.4 12.3 0.3 0.4 10.7 0.3 0.8 11.4 0.9 1.4 10.6 0.6 2.0 10.0 0.4 2.8 10.4 0.8 1.3 12.3 0.2 1.7 12.2 0.2 1.1
Guckenheimer Holmes 7.4 0.0 0.2 10.6 0.4 0.3 11.7 0.2 0.3 11.3 0.4 0.4 13.2 0.7 0.6 12.5 0.2 0.6 10.7 0.6 0.7 10.6 0.6 1.0 9.2 0.6 0.4 10.3 0.6 0.8 10.4 0.5 2.0 11.6 0.3 3.2 11.0 0.2 0.7
Hadley 8.6 0.3 0.3 10.8 0.5 0.2 9.3 0.6 0.4 11.9 0.4 0.4 12.5 0.4 0.6 12.5 0.3 0.9 10.9 0.7 0.6 10.7 0.7 2.0 9.5 0.6 0.7 10.2 0.4 1.8 11.7 0.7 1.7 12.5 0.3 5.4 11.7 0.3 0.4
Halvorsen 1.8 0.1 0.2 6.0 0.1 0.3 8.4 0.1 0.5 5.9 0.1 0.3 12.6 0.2 0.5 13.3 0.3 1.4 11.8 0.8 1.6 9.5 0.7 2.1 8.9 0.8 1.8 10.4 0.5 2.3 12.2 0.6 3.4 13.2 0.3 2.8 12.8 0.2 0.8
Hindmarsh Rose 6.1 0.5 1.1 9.7 0.3 0.3 6.7 0.5 0.7 8.4 0.1 0.3 11.5 0.4 0.9 13.8 0.6 1.5 12.7 0.7 1.5 12.5 0.7 2.2 13.9 0.6 1.2 13.4 0.7 2.2 13.9 0.6 2.5 14.0 0.2 2.7 14.2 0.6 1.5
Isothermal Chemical 2.4 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.1 13.8 0.5 0.4 7.9 0.8 0.5 5.9 0.8 0.4 8.4 0.8 0.6 8.5 0.8 0.4 12.8 0.8 0.7 15.6 0.2 6.1 4.0 0.4 0.3
Itik Banks tumor 8.7 0.2 7.0 10.0 0.6 1.7 8.4 0.5 2.4 10.5 0.4 1.5 12.8 0.3 0.8 12.1 0.7 3.2 9.2 0.7 3.6 9.5 0.7 4.6 9.7 0.8 5.1 9.6 0.7 5.1 13.3 0.9 6.4 13.2 0.3 8.2 12.7 0.3 3.8
Jerk Circuit 3.2 0.4 3.3 10.1 0.4 1.3 9.7 0.4 1.0 7.5 0.3 0.5 11.1 0.4 0.5 12.3 0.3 0.5 10.8 0.8 1.5 10.9 0.5 2.1 10.5 0.9 1.7 10.1 0.9 1.9 11.1 0.9 1.6 13.9 0.4 2.8 12.6 0.3 1.4
Laser 8.8 0.4 0.2 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.3 13.8 0.5 0.6 14.1 0.5 0.5 11.8 0.7 0.4 11.5 0.7 0.5 11.0 0.7 0.8 11.9 0.7 2.7 13.3 0.6 1.2 13.8 0.2 2.2 12.8 0.4 0.4
Lorenz 1.2 0.1 0.2 6.6 0.2 0.4 6.0 0.2 0.3 10.0 0.2 0.3 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.5 11.3 0.5 0.5 12.7 0.6 1.1 12.8 0.3 3.0 11.4 0.7 1.2 13.0 0.3 2.2 13.1 0.5 0.8
Lorenz 84 6.9 0.4 1.5 9.5 0.4 2.2 11.0 0.4 1.9 11.5 0.4 1.3 12.2 0.4 1.7 12.9 0.5 3.5 12.9 0.8 2.4 12.9 0.8 2.7 11.5 0.9 2.5 12.3 0.7 2.9 13.4 0.8 2.5 13.4 0.5 2.4 13.4 0.5 2.0
Lu Chen 7.7 0.8 0.2 6.3 0.4 0.2 5.3 0.3 0.2 8.0 0.1 0.1 7.1 0.1 0.1 4.9 0.2 0.1 11.6 0.9 0.7 11.9 1.0 0.8 11.3 1.0 0.9 11.4 1.0 0.8 9.7 0.8 0.9 13.7 0.4 1.5 8.6 0.5 0.4
Moore Spiegel 4.3 0.2 0.3 10.4 0.6 0.5 10.0 0.5 0.1 12.5 0.4 0.2 9.9 0.5 0.2 12.1 0.3 0.6 10.5 0.9 1.5 11.4 0.9 1.6 10.1 0.9 1.6 9.8 0.7 2.1 9.4 0.8 2.8 12.5 0.3 2.4 12.4 0.3 2.8
Multi Chua 8.1 0.3 0.8 9.3 0.5 0.3 9.3 0.2 0.5 7.1 0.4 0.4 13.3 0.2 1.1 14.0 0.3 3.9 11.5 0.6 5.8 12.2 0.4 11.5 9.5 0.5 11.3 10.1 0.4 9.0 10.6 0.4 4.8 14.0 0.3 7.6 11.4 0.2 2.0
Nose Hoover 9.5 0.2 0.1 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.8 13.0 0.5 0.5 10.6 0.9 0.6 10.6 0.9 0.6 10.1 0.7 0.8 10.2 0.5 1.0 10.2 0.7 1.1 12.9 0.4 1.6 13.3 0.3 0.9
Pehlivan Wei 9.2 0.0 0.2 8.5 0.4 0.5 10.1 0.4 0.5 11.3 0.3 0.3 10.8 0.4 0.5 12.6 0.5 0.7 9.2 0.7 1.2 10.1 0.8 1.7 9.2 0.7 1.0 10.6 0.8 2.1 9.3 0.7 1.4 12.2 0.6 3.0 8.2 0.3 1.2
Sakarya 2.0 0.2 0.7 5.6 0.5 0.3 5.0 0.3 0.3 7.5 0.2 0.3 11.1 0.4 0.4 12.5 0.4 0.4 10.2 0.9 1.0 10.0 0.6 1.6 7.7 0.9 2.2 9.1 0.4 1.3 13.1 0.8 1.5 12.6 0.4 1.2 11.9 0.3 2.3
San Um Srisuchinwong 3.6 0.2 0.3 12.9 0.4 1.2 12.7 0.5 2.1 11.4 0.5 1.4 13.3 0.5 2.5 13.0 0.6 1.9 12.7 0.7 1.5 10.4 0.7 2.1 13.4 0.6 1.8 11.7 0.7 1.9 10.7 0.7 2.5 13.1 0.4 3.7 13.3 0.7 1.8
Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.3 11.8 0.3 0.3 12.3 0.3 0.9 10.5 0.8 0.9 8.1 0.5 1.6 10.0 0.7 1.5 10.4 0.7 1.6 9.9 0.7 2.5 12.0 0.3 5.7 9.8 0.2 1.0
Sprott B 2.9 0.2 0.4 4.3 0.2 1.1 5.6 0.4 0.9 11.7 0.5 0.7 12.7 0.4 0.8 13.2 0.7 1.6 11.8 0.8 1.3 11.5 0.6 0.8 12.3 0.8 1.2 10.6 0.7 2.9 12.5 0.8 3.0 12.6 0.4 2.1 13.5 0.6 0.9
Sprott C 2.4 0.1 0.4 11.6 0.4 2.3 12.4 0.4 1.3 11.9 0.4 2.3 13.3 0.3 1.1 13.1 0.4 4.6 10.0 0.5 7.7 9.2 0.5 6.8 10.7 0.5 8.0 9.3 0.5 7.6 12.8 0.6 5.2 13.3 0.2 6.2 13.4 0.2 4.5
Sprott D 3.0 0.1 0.4 9.8 0.6 0.3 10.3 0.4 0.3 12.9 0.3 0.4 13.4 0.5 0.7 13.4 0.4 0.3 12.2 0.7 0.8 10.9 0.6 0.6 9.9 0.5 1.2 11.5 0.6 0.9 12.9 0.7 1.3 14.0 0.2 4.0 7.9 0.5 0.9
Sprott F 5.0 0.3 0.7 9.6 0.5 0.2 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 1.0 12.7 0.4 0.7 9.9 0.7 1.3 10.4 0.7 1.5 10.4 0.6 0.9 12.1 0.5 1.6 11.4 0.8 1.6 13.3 0.3 2.8 13.0 0.6 1.7
Sprott G 3.8 0.1 0.3 7.0 0.6 0.2 7.7 0.5 0.2 10.0 0.3 0.4 7.2 0.1 0.6 13.0 0.3 0.6 10.9 0.9 1.9 11.4 0.6 1.9 9.2 0.8 1.8 10.8 0.8 2.4 14.1 0.8 1.4 13.1 0.2 1.9 9.5 0.4 1.4
Sprott I 2.7 0.1 0.5 9.7 0.5 0.7 11.3 0.2 0.6 9.0 0.1 0.5 11.9 0.3 0.6 13.7 0.2 1.0 10.2 0.9 3.5 9.8 0.9 4.9 8.4 0.6 4.6 9.1 0.6 4.1 12.2 0.7 2.0 14.4 0.2 8.8 8.3 0.4 1.0
Sprott L 3.6 0.1 0.6 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.4 13.0 0.2 0.3 10.5 0.9 1.8 10.4 0.8 2.8 11.4 0.9 1.8 12.1 0.7 1.6 10.8 0.8 1.3 14.2 0.3 2.9 8.1 0.4 0.2
Sprott M 3.6 0.1 0.2 9.8 0.4 0.2 10.5 0.4 0.5 12.4 0.1 0.3 12.9 0.2 0.3 14.0 0.6 0.9 9.8 0.8 3.7 10.3 0.6 3.7 9.9 0.8 3.3 11.8 0.6 3.1 11.6 0.7 2.9 14.3 0.2 6.5 13.4 0.4 1.6
Sprott O 2.0 0.1 0.4 3.6 0.1 0.4 4.6 0.0 0.4 9.7 0.1 0.5 12.4 0.4 0.6 12.9 0.4 1.3 10.8 1.0 1.6 12.6 0.9 2.6 10.4 0.8 1.8 11.5 0.9 1.4 10.7 0.8 1.6 14.0 0.4 3.7 9.5 0.5 0.7
Sprott P 3.8 0.2 1.4 10.8 0.3 3.9 11.9 0.3 2.6 12.7 0.3 2.6 12.7 0.3 2.1 13.3 0.3 3.8 7.3 0.5 5.7 10.6 0.6 9.6 8.3 0.6 8.7 10.0 0.6 10.7 11.1 0.7 6.3 12.2 0.3 10.2 11.8 0.3 4.4
Sprott Q 4.5 0.1 0.2 10.1 0.2 0.3 7.6 0.1 0.4 7.3 0.1 0.5 11.6 0.2 0.4 13.2 0.3 1.3 11.8 0.8 3.6 11.6 0.7 1.9 11.9 0.6 1.0 11.2 0.6 4.1 10.9 0.7 1.1 13.1 0.4 7.5 8.8 0.4 0.4
Sprott R 4.3 0.3 0.2 9.3 0.8 0.2 6.7 0.5 0.3 11.0 0.5 0.5 12.6 0.5 0.6 13.1 0.4 0.4 10.6 0.8 0.5 10.4 0.5 1.2 9.6 0.7 2.1 10.5 0.7 1.8 11.8 0.8 1.3 13.8 0.3 2.5 11.3 0.3 0.9
Sprott S 4.2 0.1 0.3 10.1 0.9 2.2 9.9 0.6 1.1 12.3 0.2 1.2 12.6 0.4 0.6 13.1 0.3 1.5 10.6 0.7 1.2 9.7 0.8 6.1 11.0 0.6 3.1 10.4 0.4 7.2 11.8 0.8 4.1 11.6 0.3 10.2 12.8 0.2 3.7
Sprott Torus 2.2 0.2 0.2 5.4 0.7 0.4 10.4 0.7 0.3 14.3 0.6 0.3 14.5 0.6 0.6 13.8 0.7 0.4 6.2 0.8 4.2 8.3 0.6 8.3 5.3 0.8 6.3 5.8 0.7 6.3 13.6 0.8 2.4 15.5 0.3 4.8 14.3 0.5 1.0
[…]